Find the value of C such that the graph of 3x+c=5y has an intercept of (-4,0).

Answer by ankor@dixie-net.com(22740)

Find the value of C such that the graph of 3x+c=5y has an intercept of (-4,0).
:
3x + c = 5y
substitute for x and y, -4 and 0
3(-4) + c = 5(0)
-12 + c = 0
c = +12